# A $1.9\dot{9}$-approximation algorithm for the vertex cover problem on undirected, unweighted graphs(?)

### v1

@@ 6 Time Complexity Analysis
On one hand, asymptotically, $\mathcal{O}(n)$ = $\mathcal{O}(m^{2})$. This is because the maximum number of edges ($n$) possible in a simple graph is $\frac{m\cdot(m-1)}{2}$, which is less than $m^{2}$. On the other hand, asymptotically, $\mathcal{O}(n)$ = $\mathcal{O}(m)$. This is because the minimum number of edges ($n$) needed in a connected graph is $m-1$. In either case, the dominating time complexity discussed in Table 2 remains the same. In the worst case, it dominates the time complexity of all lines. In the case of a sparse graph, it either dominates or is equivalent to the time complexity of other lines. Hence, the time complexity stated in Theorem 2 holds.
@@
1 $\mathcal{O}(m\cdot\log m)$ $\mathcal{O}(m\cdot\log m)$
2 $\mathcal{O}(m^{2}n)$ $\mathcal{O}(m^{2}n)$
3 $\mathcal{O}(1)$ $\mathcal{O}(m^{2}n)$
4 $\mathcal{O}(1)$ $\mathcal{O}(m^{2}n)$
@@ 7 $\mathcal{O}(m+n)$blacktriangleright $\mathcal{O}(m^{2}+mn)$ $\mathcal{O}(m^{2}n)$
10 $\mathcal{O}(1)$blacktriangleright $\mathcal{O}(m)$ $\mathcal{O}(m^{3}n^{2})$
11 $\mathcal{O}(1)$blacktriangleright $\mathcal{O}(m)$ $\mathcal{O}(m^{3}n^{2})$
@@ 33 $\mathcal{O}(1)$ $\mathcal{O}(m^{2}n^{2})$
34 $\mathcal{O}(1)$ $\mathcal{O}(1)$
35 $\mathcal{O}(m)$ $\mathcal{O}(m)$
@@ 43 $\mathcal{O}(1)$ $\mathcal{O}(m)$ $\mathcal{O}(m^{4})$
44 $\mathcal{O}(1)$blacktriangleright $\mathcal{O}(m)$ $\mathcal{O}(m^{4})$
45 $\mathcal{O}(1)$blacktriangleright $\mathcal{O}(m)$ $\mathcal{O}(m^{4})$
46 $\mathcal{O}(m)$blacktriangleright $\mathcal{O}(m^{2})$ $\mathcal{O}(m^{4})$
47 $\mathcal{O}(m^{2})$blacktriangleright $\mathcal{O}(m^{3})$ $\mathcal{O}(m^{4})$
@@ 53 $\mathcal{O}(1)$ $\mathcal{O}(m)$ $\mathcal{O}(m^{4})$
54 $\mathcal{O}(1)$blacktriangleright $\mathcal{O}(m)$ $\mathcal{O}(m^{4})$
55 $\mathcal{O}(1)$blacktriangleright $\mathcal{O}(m)$ $\mathcal{O}(m^{4})$
56 $\mathcal{O}(1)$blacktriangleright $\mathcal{O}(m)$ $\mathcal{O}(m^{4})$
@@ 69 $\mathcal{O}(1)$ $\mathcal{O}(m^{4})$
70 $\mathcal{O}(m+m)$ $\mathcal{O}(m)$
71 $\mathcal{O}(m)$ $\mathcal{O}(m)$
72 $\mathcal{O}(1)$ $\mathcal{O}(m)$
73 $\mathcal{O}(m^{2})$ $\mathcal{O}(m^{2})$
74 $\mathcal{O}(m+m)$ $\mathcal{O}(m^{2})$
@@ 7 Conclusion
We show that the VC problem can be solved efficiently. It implies that DiRe committees can be computed efficiently. Hence, achieving diversity and representation is more efficient than initially expected. Also, indeed, P $=$ NP.
Broader Impact:
We do not expect major, immediate, positive or negative, practical implications of this work. It is primarily because extrapolating our algorithm to elections where candidates are divided into arbitrarily sized arbitrary groups itself seems non-trivial (a.k.a. extrapolating our algorithm to hypergraphs itself seems non-trivial).
@@ Appendix C Vertex Cover on Unweighted Simple Connected Graphs
    vertex cover $S^{\prime}$ consists of zero dummy vertex from $D^{\prime}$, $k$ vertices from $V^{\prime}$, one vertex $u^{\prime}$ This is a trivial case and the instance of the VC1 problem will have vertex cover $S$ consisting of vertex $v_{i}$ for every vertex $v^{\prime}_{i}\in S^{\prime}$. This will be of size $k$.
@@
    vertex cover $S^{\prime}$ consists of $x$141414variable $x$ is an integer such that $1\leq x\leq k$. dummy vertex from $D^{\prime}$, $k-x$ vertices from $V^{\prime}$, one vertex $u^{\prime}$ For each dummy vertex $d^{\prime}\in D^{\prime}$ selected, the corresponding vertex $v^{\prime}_{i}\in V^{\prime}$ connected to the dummy vertex is not selected. Hence, given that the dummy vertex is of degree one, it can be swapped with the vertex it is connected to. This won’t have any effect on the validity of the vertex cover $S^{\prime}$. In summary, $x$ dummy vertices from $D^{\prime}$ in vertex cover $S^{\prime}$ are replaced by the corresponding $x$ vertices from $V^{\prime}$. Consequently, an instance of the VC1 problem will have vertex cover $S$ consisting of vertex $v_{i}$ for every vertex $v^{\prime}_{i}\in S^{\prime}$. This will be of size $k$.
@@
    vertex cover $S^{\prime}$ consists of zero dummy vertex from $D^{\prime}$, zero vertices from $V^{\prime}$ and one vertex $u^{\prime}$ In such a case, one endpoint of all edges in VC2 is $u^{\prime}$. Hence, the corresponding instance of VC1 contains no edges and its vertex cover will be a null set.
Finally, note that no other cases can lead to a yes instance of VC2 (e.g., $S^{\prime}$ is not a vertex cover if $S^{\prime}$ consists of, for example, $x$ dummy vertices from $D^{\prime}$, $k-x+1$ vertices from $V^{\prime}$ and zero vertex from $U^{\prime}$).

### v2

@@ 5 Time Complexity Analysis
On one hand, asymptotically, $\mathcal{O}(n)$ = $\mathcal{O}(m^{2})$. This is because the maximum number of edges ($n$) possible in a simple graph is $\frac{m\cdot(m-1)}{2}$, which is less than $m^{2}$. On the other hand, asymptotically, $\mathcal{O}(n)$ = $\mathcal{O}(m)$. This is because the minimum number of edges ($n$) needed in a connected graph is m1m-1. In either case, the dominating time complexity discussed in Table 2 remains the same. In the worst case, it dominates the time complexity of all lines. In the case of a sparse graph, it either dominates or is equivalent to the time complexity of other lines. Hence, the time complexity stated in Theorem 1 holds.
@@
1 $\mathcal{O}(m\cdot\log m)$ - $\mathcal{O}(m\cdot\log m)$
2 $\mathcal{O}(m^{2}n)$ - $\mathcal{O}(m^{2}n)$
3 $\mathcal{O}(1)$ - $\mathcal{O}(m^{2}n)$
4 $\mathcal{O}(1)$ - $\mathcal{O}(m^{2}n)$
@@ 6 $\mathcal{O}(1)$ $\mathcal{O}(m)$ $\mathcal{O}(m^{2}n)$
[…]
@@ 33 $\mathcal{O}(1)$ - $\mathcal{O}(m^{2}n^{2})$
34 $\mathcal{O}(1)$ - $\mathcal{O}(1)$
35 $\mathcal{O}(m)$ - $\mathcal{O}(m)$
@@ 43 $\mathcal{O}(1)$ $\mathcal{O}(m)$ $\mathcal{O}(m^{4})$
[…]
@@ 69 $\mathcal{O}(1)$ - $\mathcal{O}(m^{4})$
70 $\mathcal{O}(m+m)$ - $\mathcal{O}(m)$
71 $\mathcal{O}(m)$ - $\mathcal{O}(m)$
72 $\mathcal{O}(1)$ - $\mathcal{O}(m)$
73 $\mathcal{O}(m^{2})$ - $\mathcal{O}(m^{2})$
74 $\mathcal{O}(m+m)$ - $\mathcal{O}(m^{2})$
@@ 6 Conclusion
We conjecture that the VC problem can be solved efficiently. It implies that DiRe committees can be computed efficiently. Hence, achieving diversity and representation may be more efficient than initially expected.
@@ Appendix C Vertex Cover on Unweighted Simple Connected Graphs
    vertex cover $S^{\prime}$ consists of zero dummy vertex from $D^{\prime}$, $k$ vertices from $V^{\prime}$, one vertex $u^{\prime}$ - This is a trivial case and the instance of the VC1 problem will have vertex cover $S$ consisting of vertex $v_{i}$ for every vertex $v^{\prime}_{i}\in S^{\prime}$. This will be of size $k$.
@@
    vertex cover $S^{\prime}$ consists of xx101010variable $x$ is an integer such that $1\leq x\leq k$. dummy vertex from $D^{\prime}$, kxk-x vertices from $V^{\prime}$, one vertex $u^{\prime}$ - For each dummy vertex $d^{\prime}\in D^{\prime}$ selected, the corresponding vertex $v^{\prime}_{i}\in V^{\prime}$ connected to the dummy vertex is not selected. Hence, given that the dummy vertex is of degree one, it can be swapped with the vertex it is connected to. This won’t have any effect on the validity of the vertex cover $S^{\prime}$. In summary, $x$ dummy vertices from $D^{\prime}$ in vertex cover $S^{\prime}$ are replaced by the corresponding $x$ vertices from $V^{\prime}$. Consequently, an instance of the VC1 problem will have vertex cover $S$ consisting of vertex $v_{i}$ for every vertex $v^{\prime}_{i}\in S^{\prime}$. This will be of size $k$.
@@
    vertex cover $S^{\prime}$ consists of zero dummy vertex from $D^{\prime}$, zero vertices from $V^{\prime}$ and one vertex $u^{\prime}$ - In such a case, one endpoint of all edges in VC2 is $u^{\prime}$. Hence, the corresponding instance of VC1 contains no edges and its vertex cover will be a null set.
Finally, note that no other cases can lead to a yes instance of VC2 (e.g., $S^{\prime}$ is not a vertex cover if $S^{\prime}$ consists of, for example, $x$ dummy vertices from $D^{\prime}$, kx+1k-x+1 vertices from $V^{\prime}$ and zero vertex from $U^{\prime}$).